OpenStudy (unklerhaukus):

Solve the ODE \[\frac{\text dy}{\text dx}=\frac{2y-x-1}{4x-3y-6}\]

OpenStudy (unklerhaukus):

i got \[(3y+x-9)^5=A(y-x+1)\] how can i check it?

OpenStudy (unklerhaukus):

where A is a constant

OpenStudy (unklerhaukus):

so to check if it is a solution to the Differential Equation, we differentiate right? and hopeful get back the original DE,

OpenStudy (unklerhaukus):

my working

OpenStudy (unklerhaukus):

???

OpenStudy (unklerhaukus):

also if you can show me a was to solve the DE that dosent take three pages that would be helpful too

OpenStudy (apoorvk):

Well your solution seems perfectly alright to me - I went through it step by step, and it seems okay - except where you differentiate it - am not sure if that's okay too (the last too steps). This problem can be shortened (according to my knowledge) only at the point where you have used partial fractions - instead of that why dont you make a whole square in the denominator, and then split the numerator?

OpenStudy (apoorvk):

**two steps

OpenStudy (anonymous):

sorry unkle.. anything with more 1/2 a page of work i start seeing pink elephants...

OpenStudy (unklerhaukus):

this is just a single problem from a set of 12 i have been working on the set all year

OpenStudy (unklerhaukus):

This is that last one

OpenStudy (apoorvk):

Unkle can you make whole square of the rm instead of using partial fractions??

OpenStudy (unklerhaukus):

i am not sure how to do that do you mean after this step?\[\frac{\text dX}{X}=\frac{4-3V}{3V^2-2V-1}\text dV\]

OpenStudy (apoorvk):

yes.

OpenStudy (unklerhaukus):

\frac{\text dX}{X}=\frac{4-3V}{3V^2-2V-1}\text dV

OpenStudy (apoorvk):

You know what - what you did was the best possible method actually - creating whole squares is pretty tedious in the current arrangement of numbers. - I guess this is the only way you'll have to proceed in this question.

OpenStudy (unklerhaukus):

i am not sure what the whole square method is is the same as completing the square? can you provide a simple example

OpenStudy (anonymous):